# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ Abstract
journal: some journal
@@
Department of Foundations of Computer Science University of Szeged, addressline=Arpad ter 2, city=Szeged, postcode=H-6720, country=Hungary
{highlights}
We can decide in square time whether the union of two finitely generated congruences on the ground term algebra is also a congruence.
If the answer is ’yes’, then the union of the two generator sets of ground equations generates the union of the congruences.
@@ 1 Introduction
Vágvölgyi [1] showed that it is decidable for any given ground term equation systems (GTESs for short) $E$ and $F$ whether there is a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cap\mathop{\leftrightarrow}^{*}_{F}$. If the answer is yes, then we can construct such a GTES $H$. We show the dual to the above result, first we show that for any GTESs $E$ and $F$, $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra if and only if there exists a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Second, we show that it is decidable in $O(n^{2})$ time for any given GTESs $E$ and $F$ whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n$ is the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$.
@@
For each signature $\Sigma$ without constants, the set of ground terms over $\Sigma$ is the empty set, each GTES over $\Sigma$ is the empty set, and our results above are obvious. Thus, throughout the paper we consider signatures with constants. We now sketch the proof of our main result. Let $E$ and $F$ be GTESs over $\Sigma$, and let $W$ denote the set of subterms appearing in the GTESs $E$ and $F$. Fülöp and Vágvölgyi [2] presented a fast ground completion algorithm which, given GTESs $E$ and $F$ over a signature $\Sigma$, produces in $O(n\,\mathrm{log}\;n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced ground term rewrite system (GTRS for short) $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Here $n$ stands for the number of occurrences of symbols in $E$ plus number of occurrences of symbols in $F$. Furthermore, the set $C\langle E,F\rangle$ of new constants is the set of the equivalence classes of the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to the set $W$. For this purpose we proceed as follows:
[…]
We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
Then we decide in $O(n)$ time whether $R\langle E,F\rangle$ and $R\langle F,E\rangle$ are total. Accordingly, we distinguish the following four main cases.
@@
Main Case 1: $\Sigma$ is a unary signature. In this main case $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Condition (i) or Condition (ii) hold:
  • (i)
    • (a)
      $a\in C\langle E,F\rangle$ and
    • (b)
      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]\rangle=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E,F\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
  • (ii)
    • (a)
      $a\in C\langle F,E\rangle$ and
    • (b)
      for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
We decide the above formulas by carrying out a nested partial depth first search on the auxiliary dpwpa $AUX[E;F]$. In the outer loop we carry out a partial depth first search on $AUX[E;F]$ and decide Conditions (i).(a) and (ii).(a). In the inner loop we visit those vertices which are positive-step reachable from the current vertex of the outer loop, and examine at each visited vertex $a$ the value of the Boolean attribute $keeps_{E}$ (respectively $keeps_{F}$), which tells us whether $R\langle E,F\rangle$ (resp. $R\langle F,E\rangle$) keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $a$. In this way we decide Condition (ii).(a) (resp. Condition (ii).(b)). In such manner, we decide in additional $O(n^{2})$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
Main Case 2: Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class. We decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
@@
Main Case 3: $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total. Without loss of generality we may assume that $R\langle E,F\rangle$ is total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$,
  • (1)
    $a\in C\langle E,F\rangle$ or $a\in C\langle F,E\rangle$ and
  • (2)
    if $a\not\in C\langle E,F\rangle$, then for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle F,E\rangle$ keeps up with $R\langle E\cup F,\emptyset\rangle$ writing $b$.
In $O(n)$ time, we decide set memberships $a\in C\langle E,F\rangle$ and $a\in C\langle F,E\rangle$ for all $a\in C\langle E\cup F,\emptyset\rangle$. Therefore we decide in $O(n)$ time whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$. We decide in $O(n)$ time whether Condition (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ by carrying out a partial depth first search on the auxiliary dpwpa $AUX[E;F]$ by narrowing down the set of vertices from which we start a session of the search. We start the search sessions from those $C\langle E\cup F,\emptyset\rangle$ vertices $a$ which are not in $C\langle E,F\rangle$ and have not been visited previously. If we start a search session from a vertex $a$ and there exists a cycle leading from the vertex $a$ to itself, then along this cycle Algorithm PDFS visits vertex $a$ for a second time. During the current session of the search we find a vertex $b$ with false value for the attribute $keeps_{F}$ if and only if Condition (2) does not hold for the vertex $a\in C\langle E\cup F,\emptyset\rangle$ from which we started the current session of the search. We decide whether Condition (1) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ and whether Conditions (2) holds for all $a\in C\langle E\cup F,\emptyset\rangle$ in $O(n)$ time, in this way, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Main Case 4: $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$ if and only if $E\subseteq\Theta\langle F,E\rangle$ or $F\subseteq\Theta\langle E,F\rangle$. In this case, by direct inspection of the representations $\tau\langle E,F\rangle$ of $E$, $\tau\langle F,E\rangle$ of $F$, $\rho\langle E,F\rangle$ of ${\Theta\langle E,F\rangle}$ and $\rho\langle F,E\rangle$ of ${\Theta\langle F,E\rangle}$ in the subterm dag for $E$ and $F$, we decide in additional $O(n)$ time whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Summing up the previous thoughts, our decision algorithm goes as follows: We compute the following sets and relations and dpwpa in $O(n\,\mathrm{log}\,n)$ time:
  • the GTRS $R\langle E\cup F,\emptyset\rangle$ stored in a red-black tree;
  • the GTRS $R\langle E,F\rangle$ stored in a red-black tree;
  • the GTRS $R\langle F,E\rangle$ stored in a red-black tree;
  • the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$ for the GTESs $E$ and $F$, and for every vertex $a\in C\langle E\cup F,\emptyset\rangle$, the values of the attributes $a.equal_{E}$, $a.equal_{F}$, $a.keeps_{E}$, and $a.keeps_{F}$.
We decide whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total.
  • If both of them are total, then we proceed as in main Case 2.
  • Otherwise,
    • if $\Sigma$ is a unary alphabet, then we procced as in Main Case 1.
    • If $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total, the we proceed as in Main Case 3.
@@
    • If $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total, then we proceed as in Main Case 4.

### v2

@@ Abstract
journal: Journal of Automata, Languages and Combinatorics
@@
Department of Foundations of Computer Science University of Szeged, addressline=Árpád tér 2, city=Szeged, postcode=H-6720, country=Hungary
@@ 1 Introduction
Kozen [1], Nelson and Oppen [2], and Tarjan [3] studied the word problem for ground term equation systems (GTESs) and applied the graph-based method of congruence closure. In order to decide for a given GTES $E$ and ground terms $s$ and $t$, whether $s\mathop{\leftrightarrow}^{*}_{E}t$ holds, they constructed the subterm graph of $E$ and $s$ and $t$ and computed the congruential consequences of $E$ on the vertices of the graph. Kozen [1] showed that the word problem for every GTES is decidable in polynomial time. Nelson and Oppen [2] and Tarjan [3] showed that the word problem for every GTES is decidable in $O(n\,\mathrm{log}\,n)$ time applying the congruence closure procedure on the subterm graph of the GTES $E$ and ground terms $s$ and $t$. Continuing this research line, Snyder [4] gave an efficient algorithm which, given a GTES $E$ over a signature $\Sigma$, produces an equivalent, reduced ground term rewriting system $R$ over $\Sigma$ in $O(n\,\mathrm{log}\,n)$ time. Then Fülöp and Vágvölgyi [5] gave a simple and efficient algorithm which, given a GTES $E$ over a signature $\Sigma$, produces a reduced ground term rewriting system (GTRS for short) $R$ over a signature containing $\Sigma$ such that two terms are congruent modulo $E$ if and only if their $R$-normal forms are equal. The algorithm calls a congruence closure procedure [3] on the subterm graph of the GTES $E$ and runs in $O(n\,\mathrm{log}\,n)$ time, where $n$ is the size of $E$. We say that a tree language is congruential if it is the union of finitely many congruence classes of a finitely generated congruence. Brainerd [6], Kozen [1] showed that congruential tree languages are the same as recognizable tree languages. Applying the above algorithm of Fülöp and Vágvölgyi [5], for a given GTES $E$ and ground terms $p_{1},\ldots,p_{k}$ of size $n$, a deterministic bottom-up tree automaton can be constructed in $O(n\,\mathrm{log}\,n)$ time which recognizes the congruential tree language $p_{1}/_{\mathop{\leftrightarrow}^{*}_{E}}\cup\cdots\cup p_{k}/_{\mathop{\leftrightarrow}^{*}_{E}}$ [7].
@@
Vágvölgyi [8] showed that it is decidable for any given GTESs $E$ and $F$ whether there is a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cap\mathop{\leftrightarrow}^{*}_{F}$. If the answer is “yes”, then we can construct such a GTES $H$. We show the dual to the above result, first we show that for any GTESs $E$ and $F$, $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ is a congruence on the ground term algebra if and only if there exists a GTES $H$ such that $\mathop{\leftrightarrow}^{*}_{H}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Second, we show that it is decidable in $O(n^{2})$ time for any given GTESs $E$ and $F$ whether $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, where $n$ is the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$. For each signature $\Sigma$ without constants, the set of ground terms over $\Sigma$ is the empty set, each GTES over $\Sigma$ is the empty set, and our results above are obvious. Thus, throughout the paper we consider signatures with constants.
@@
We now recall the fast ground completion algorithm of Fülöp and Vágvölgyi [5]. Let $E$ and $F$ be GTESs over $\Sigma$, and let $W$ denote the set of subterms appearing in the GTESs $E$ and $F$. Fülöp and Vágvölgyi [5] presented a fast ground completion algorithm which, given GTESs $E$ and $F$ over a signature $\Sigma$, produces in $O(n\,\mathrm{log}\,n)$ time a finite set $C\langle E,F\rangle$ of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Here $n$ stands for the number of occurrences of symbols in $E$ plus the number of occurrences of symbols in $F$. Furthermore, the set $C\langle E,F\rangle$ of new constants is the set of the equivalence classes of the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to the set $W$. For this purpose we proceed as follows: In the first step in $O(n)$ time, we create a subterm directed, labeled, and acyclic multigraph (dag) $G$ for $E$ and $F$ [4]. We consider the vertices in the subterm dag $G$ as terms in $W$. The relation $\tau\langle E,F\rangle$ on the vertices of $G$ corresponds to $E$. We construct the relation $\tau\langle E,F\rangle$. Then we run on $G$ a congruence closure algorithm for the relation $\tau\langle E,F\rangle$ in $O(n\,\mathrm{log}\,n)$ time [9]. In this way, we obtain on $G$ the congruence closure $\rho\langle E,F\rangle$ of the relation $\tau\langle E,F\rangle$. Considering the vertices in the subterm dag $G$ as terms, $\rho\langle E,F\rangle$ yields the restriction $\Theta\langle E,F\rangle$ of $\mathop{\leftrightarrow}^{*}_{E}$ to $W$. Then using $\rho\langle E,F\rangle$, $\Theta\langle E,F\rangle$, and $C\langle E,F\rangle$, we produce the GTRS $R\langle E,F\rangle$ in $O(n)$ time. GTRS $R\langle E,F\rangle$ is the set of all rewrite rules $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$, where $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m},{b}\in C\langle E,F\rangle$, and $\{\,\sigma(t_{1},\ldots,t_{m})\mid t_{1}\in{b}_{1},\ldots,t_{m}\in{b}_{m}\,\}\subseteq{b}$. Fülöp and Vágvölgyi [5] showed that GTRS $R\langle E,F\rangle$ is reduced, and hence is convergent. We construct GTRSs $R\langle F,E\rangle$ and $R\langle E\cup F,\emptyset\rangle$ in a similar way.
The intuitive idea of the paper is as follows. Let $E$ and $F$ be GTESs over $\Sigma$. For every subterm $t$ appearing in $E\cup F$, GTRS $R\langle E,F\rangle$ reduces all elements of the congruence class $t/_{\mathop{\leftrightarrow}^{*}_{\langle E,F\rangle}}$ to the constant (equivalence class) $t/_{\Theta\langle E,F\rangle}$, and GTRS $R\langle E\cup F,\emptyset\rangle$ reduces every element of the congruence class $t/_{\mathop{\leftrightarrow}^{*}_{\langle E\cup F,\emptyset\rangle}}$ to the constant (equivalence class) $t/_{\Theta\langle E\cup F,\emptyset\rangle}$. For every subterm $t$ appearing in $E\cup F$, the congruence class $t/_{{\mathop{\leftrightarrow}^{*}_{\langle E\cup F,\emptyset\rangle}}}$ is equal to the congruence class $t/_{{\mathop{\leftrightarrow}^{*}_{\langle E,F\rangle}}}$ if and only if $t/_{\Theta\langle E\cup F,\emptyset\rangle}=t/_{\Theta\langle E,F\rangle}$ and for every $p\in T_{\Sigma}$ and reduction sequence ${\cal RS}1$ from $p$ to $t/_{\Theta\langle E\cup F,\emptyset\rangle}$ by $R\langle{E\cup F},\emptyset\rangle$, there exists a reduction sequence ${\cal RS}2$ from $p$ to $t/_{\Theta\langle E,F\rangle}$ by $R\langle E,F\rangle$, where each rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ applied at any node of $p$ along the reduction sequence ${\cal RS}1$ is matched with some rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ in $R\langle E,F\rangle$ applied by GTRS $R\langle E,F\rangle$ at the same node along the reduction sequence ${\cal RS}2$. By matching we mean that ${c}_{i}\subseteq{b}_{i}$ for all $i\in\{\,1,\ldots,m\,\}$, and ${c}\subseteq{b}$. For every ${b}\in C\langle E\cup F,\emptyset\rangle$, we say that GTRS $R\langle E,F\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b} if the following holds. If GTRS $R\langle E\cup F,\emptyset\rangle$ applies a rewrite rule of the form $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ along a reductions sequence from $p$ to $t/_{\Theta\langle E,F\rangle}$, then $R\langle E,F\rangle$ also applies a rewrite rule $\sigma({c}_{1},\ldots,{c}_{m})\dot{\rightarrow}{c}$ for some ${c}_{1},\ldots,{c}_{m},c\in C\langle E,F\rangle$ at the same node, that is, $R\langle E,F\rangle$ also move forward along its reduction sequence. For every ${b}\in C\langle E\cup F,\emptyset\rangle$, we define that GTRS $R\langle F,E\rangle$ keeps up with the GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b} in a similar way.
For each constant ${d}\in C\langle E\cup F,\emptyset\rangle$, we keep track of all constants ${c}\in C\langle E\cup F,\emptyset\rangle$ which appear as a right-hand side of a rewrite rule applied along the reduction sequence from some ground term $t\in T_{\Sigma}$ to the constant d{d} by $R\langle E\cup F,\emptyset\rangle$. To this end, we construct a directed pseudograph without parallel arcs (dpwpa), $AUX[E;F]$ with vertices $C\langle E\cup F,\emptyset\rangle$. At each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$, we indicate whether it is an element of $C\langle E,F\rangle$ whether it is an element of $C\langle F,E\rangle$, whether GTRS $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b}, and whether GTRS $R\langle F,E\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing the constant b{b}. Every directed arc (b,c)({b},{c}) is associated with some rewrite rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}$ in $R\langle E\cup F,\emptyset\rangle$, where ${c}={b}_{i}$ for some $i\in\{\,1,\ldots,m\,\}$. That is, b{b} is the right-hand side of the rewrite rule, and c{c} appears on the left-hand side of the rewrite rule. We say that GTRS $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$ if ${d}\in C\langle E,F\rangle$ and for all constants c{c} appearing in a reduction sequence from a ground term $p\in T_{\Sigma}$ to d{d} by GTRS $R\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $c$. That is, for all constants c{c} with a path of positive length or a cycle from c{c} to d{d} in $AUX[E;F]$, $R\langle E,F\rangle$ keeps up with GTRS $R\langle E\cup F,\emptyset\rangle$ writing $c$. We define that GTRS $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$ in a similar way. During the run of the decision algorithm, we carry out a partial depth first search on the auxiliary dpwpa $AUX[E;F]$ to decide whether $R\langle E,F\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$, and to decide whether $R\langle F,E\rangle$ simulates the GTRS $R\langle E\cup F,\emptyset\rangle$ on a constant ${d}\in C\langle E\cup F,\emptyset\rangle$. We say that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ if for each constant ${d}\in C\langle E\cup F,\emptyset\rangle$, GTRS $R\langle E,F\rangle$ or GTRS $R\langle F,E\rangle$ simulate the GTRS $R\langle E\cup F,\emptyset\rangle$ on d{d}. We show that for every subterm $t$ appearing in $E\cup F$, $\left(t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{E}}}\mbox{ or }t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{{\mathop{\leftrightarrow}^{*}_{F}}}\right)$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. We consider the following four main cases.
@@
Main Case 1: $\Sigma$ is a unary signature. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
Main Case 2: Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if each $\Theta\langle E\cup F,\emptyset\rangle$ equivalence class is equal to a $\Theta\langle E,F\rangle$ equivalence class or a $\Theta\langle F,E\rangle$ equivalence class.
@@
Main Case 3: $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$ and $R\langle E,F\rangle$ or $R\langle F,E\rangle$ is total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
Main Case 4: $\Sigma$ has a symbol of arity at least $2$, and GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are not total. Then $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ if and only if $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$ or $F\subseteq\mathop{\leftrightarrow}^{*}_{E}$.